Books and Brew (BB) is a large city bookstore that sells books and music CDs, and also has a cafe. Currently, BB uses a single-driver system to allocate its operating costs to each of its three product lines, using the number of items sold as the single cost driver. But BB's management is concerned that this allocation system may not be providing the best information for making a variety of pricing decisions. BB's operating costs for 2019 were as follows:

Purchasing department $500,000
Receiving department   $449,000
Shelf-stocking employee salaries   $483,000
Cashiers and floor employee salaries   $147,000

2019 information about BB's product lines is also available:

Books CDs Cafe
Revenue $3,921,000 $2,390,000 $787,000
Cost of merchandise $2,690,000 $1,681,000 $575,000
Number of purchase orders placed 2,660 2,570 2,000
Number of deliveries received 1,480 1,500 1,490
Hours of shelf stocking time 15,200 14,100 10,300
Number of items sold 112,000 107,000 272,000


REQUIRED [ROUND ALLOCATION RATES TO THE NEAREST CENT AND ALLOCATIONS TO THE NEAREST DOLLAR.]

Part A [6 tries; 5 points]
Using BB's single-driver system to allocate its operating costs, how much was allocated to CDs in 2019?

You have entered that answer before Incorrect. Tries 3/6 Previous Tries


Part B [6 tries; 5 points]
If BB had used an activity-based costing system to allocate its operating costs in 2019, how much would have been allocated to Books? For cashiers and floor employee salaries, use number of items sold as the cost driver; for the other three cost categories, it should be clear what the appropriate drivers are.
Answer #1
Part-A Total Operating Cost $ 1,579,000 =500000+449000+483000+147000
Total Number of units sold    491,000 =112000+107000+272000
Allocation rate $ 3.22 =1579000/491000
Operating costs allocated to CDs in 2019 $ 344,540 =107000*3.22
Part-B A B A/B
Cost $ Total Driver Base Allocation rate Allocation Base
Purchasing department 500,000                7,230                  69.16 Per number of purchase order placed
Receiving department 449,000                4,470               100.45 Per number of deliveries received
Shelf-stocking employee salaries 483,000              39,600                  12.20 Per hour of shelf stocking time
Cashiers and floor employee salaries 147,000          491,000                    0.30 Per number of item sold
Amount $
Purchasing department 177,741 =2570*69.16
Receiving department 150,675 =1500*100.45
Shelf-stocking employee salaries 172,020 =14100*12.2
Cashiers and floor employee salaries 32,100 =107000*0.3
Operating costs allocated to books 532,536
Working:
A B A/B
Cost $ Driver Base Allocation rate Allocation Base
Purchasing department 500,000 =2660+2570+2000                  69.16 Per number of purchase order placed
Receiving department 449,000 =1480+1500+1490               100.45 Per number of deliveries received
Shelf-stocking employee salaries 483,000 =15200+14100+10300                  12.20 Per hour of shelf stocking time
Cashiers and floor employee salaries 147,000 =112000+107000+272000                    0.30 Per number of item sold